PUBLICATIONS OF J.B. CONREY

1. (with L.A. Rubel) On the location of the zeros of the derivatives of a polynomial, Proc. A.M.S., 86 (1982), 37–41.

2. Zeros of derivatives of Riemann’s xi-function on the critical line, J. Number Th., 16 (1983), 49–74.

3. Zeros of derivatives of Riemann’s xi-function on the critical line, J. Number Th., 17 (1983), 71–75.

4. (with A. Ghosh and S.M. Gonek) A note on gaps between zeros of the zeta-function, Bull. London Math. Soc., 16 (1984), 421–424.

5. (with A. Ghosh) On mean values of the zeta-function, Mathematika, 31 (1984), 159–161.

6. (with A. Ghosh, D. Goldston, S.M. Gonek and D.R. Heath-Brown) Distribution of gaps between zeros of the zeta-function, Quarterly J. of Math. Oxford (2), 36 (1985), 43–51.

7. (with R. Balasubramanian, and D.R. Heath-Brown) Asymptotic mean square of the product of the Riemann zeta-function and a Dirichlet polynomial, J. Reine Agnew. Math., 357 (1985), 161–181.

8. (with A. Ghosh) A simpler proof of Levinson’s theorem, Math. Proc. Cambridge Phil. Soc., 97 (1985), 385–395.

9. On the distribution of the zeros of the Riemann zeta-function, \underline{Topics in Analytic} \underline{Number Theory}, University of Texas Press, Austin, Texas, 1985.

10. (with A. Ghosh) A mean value theorem for the Riemann zeta-function at its relative extrema on the critical line, J. London Math. Soc. (2) 32 (1985) 193–202.

11. (with A. Ghosh and S.M. Gonek) Large gaps between zeros of the zeta-functions, Mathematika, 33 (1986), 212–238.

12. (with A. Ghosh and S.M. Gonek) Simple zeros of the zeta-function of a quadratic number field, I, Inventiones Math., 86 (1986), 536–576.

13. (with A. Ghosh and S.M. Gonek) Simple zeros of the zeta-function of a quadratic number field, II, \underline{Analytic Number Theory and Diophantine Problems}, Birkhäuser, Boston · Basel · Stuttgart, 1987.

14. The fourth moment of derivatives of the Riemann zeta-function, Quarterly J. Math. Oxford, (2) 39 (1988), 21–36.

15. (with A. Ghosh) On the zeros of the Taylor polynomials associated with the exponential function, American Math. Monthly, 95 (1988), 528–533.

16. (with A. Ghosh) Simple zeros of the Dirichlet series formed with Ramanujań tau–function, C. R. Math. Rep. Acad. Sci. Canada, 10, (1988), 95–99.

17. (with A. Ghosh) Simple zeros of the Ramanujan τ-Dirichlet series, Inventiones Math., 94 (1988), 403–419.

18. At least two-fifths of the zeros of the Riemann zeta-function are on the critical line, Bull. AMS., 20 (1989), 79-81.

19. (with A. Ghosh and S.M. Gonek) Mean values of the Riemann zeta-function with application to the distribution of zeros, in Number Theory, Trace Formulas and Discrete Groups, Symposium in Honor of Atle Selberg, Academic Press, San Diego, 1989.

20. More than two-fifths of the zeros of the Riemann zeta-function are on the critical line, J. Reine Angew. Math., 399 (1989), 1-26.

21. (with A. Ghosh) On mean values of the zeta-function, II, Acta Arithmetica, LII (1989), 367-371.

22. (with Michael W. Smiley) Some geometric aspects of hyperbolic boundary value problems, Proc. American Math. Soc., 107 (1989), 591-601.

23. (with A. Ghosh) Zeros of derivatives of the Riemann zeta-function, \underline{Analytic Number Theory}, Proceedings of a Conference in Honor of Paul J. Bateman, Birkhäuser, Boston· Basel· Berlin(1990), 95-110.

24. (with A. Ghosh) Mean values of the Riemann zeta-function, III, in Proceedings of the Amalfi Conference on Analytic Number Theory, Università di Salerno, 1992, 35-59.

25. (with A. Ghosh) On the Selberg class of Dirichlet series: small degrees, Duke, 72 (1993), 673-693.

26. (with A. Ghosh) Turan inequalities and zeros of Dirichlet series associated with certain Dirichlet series. Trans. American Math. Soc., 342 (1994), 407-419.

27. (with D. Farmer) An extension of Hecke’s converse theorem, International Mathematics Research Notices, 9 (1995) 445-463.

28. (with E. Fransen, R. Klein, C. Scott) Mean Values of Dedekind Sums, J. Number Theory, 56 (1996) 214 - 226.

29. A note on the fourth power moment of the Riemann zeta-function, Analytic Number Theory: Proceedings of a Conference in Honor of Heini Halberstam, Volume 1, Birkhauser, 1996, 225 - 230.

30. (with P. M. Agrawal, D. C. Sorescu, R. D. Kay, D. L. Thompson, L. M. Raff, and A. K. Jameson) Spectral lines shapes in systems undergoing continuous frequency modulation, Journal of Chemical Physics, 105 (1996) 2686-2700.

PREPRINTS:

31. (with D. Farmer) The distribution of the eigenvalues of Hecke operators, submitted to Acta Arithmetica.

32. (with A. Ghosh and S. M. Gonek) Simple zeros of the Riemann zeta-function, submitted to JLMS.

33. (with A. Ghosh) Remarks on the Generalized Lindelöf Hypothesis.

34. (with A. Granville and B. Poonen) Zeros of Fekete Polynomials. BOOK:

35. (co–edited with A. Adolphson, A. Ghosh, and R. Yager) Analytic Number Theory and Diophantine Problems, Birkhäuser, Boston · Basel · Stuttgart, 1987.